Physics :

A proton (Charge= +e, mass=M) and a deuteron (charge = +e, mass = 2M) have the same kinnetic energy. They enter a magnetic field and both move in circular paths. How are the radii of the paths Rp and Rd related?

A.) Rd = Rp

B.) Rd = 2Rp

C.). Rd = sqrt(2Rp)

D.) Rd = Rp/2

E.) Rd = Rp / sqrt(2)
Answer #1


E) Rd = Rp / sqrt(2)

we know,

radius of orbit, r = m*v/(B*q)

= sqrt(2*m*KE)/(B*q)


let KE is kinetic energy of the two particles.
so,

Rp = sqrt(2*M*KE)/(B*e)


Rd = sqrt(2*2*M*KE)/(B*2*e)


take the ration,

Rd/Rp = ( sqrt(2*2*M*KE)/(B*2*e) )/(sqrt(2*M*KE)/(B*e) )

Rd/RP = sqrt(2)/2

Rd/RP = 1/sqrt(2)

Rd = Rp/sqrt(2)